We used the problem solver for this equation y/2 +3=5
I am still very confused so it says C
move the 5 to the left it becomes y/2 +3-= WHY DID WE MOVE THE 5 TO THE LEFT? HOW DID WE KNOW TO DO THAT? AND WHY THE 5 OUT OF ALL THE NUMBERS IN THE EQUATION?
add fractions or integers together it becomes y/2+-2=0 I DONT UNDERSTAND ANY OF THIS.
remove extra sign in front of -2 y/2-2=0
solved y/2-2=0
it becomes 0
this equation y=4
WOW why isnt it 0? I'm completely lost, maybe you can somehow explain the order and reason for each step, please
Thank you so much for your time.

Answer by Boreal(15235)

Here is how I would deal with
y/2 +3= 5. Then I will show you how they dealt with it.
I like to separate the numbers and the variables.
I wold move the 3 to the right. Then all the numbers are on the right, and the y is on the left. y/2 is considered part of y.
To move the 3 to the right, I subtract 3 from both sides
y/2=2, because 5-3 on the right is 2.
Now, to get rid of the denominator, I have (1/2) y or y/2, they are the same, I multiply both sides by 2.
2*y/2=y. I now have y by itself and just 1y. That is what I want. But I have to multiply the other side by 2
y=4, because 2*2=4.
---------------------------
You can move the 5 to the left
then it is y/2-2=0
but then I need to move the 2 to the right by adding 2
I get y/2=2
Then I do what I did before.
It isn't wrong, but to me it is another step I don't need to do.
I can also take y/2-2=0 and multiply both sides by 2
y/2 *2=y, -2*2=-4, and 0*2=0
y-4=0
Add 4 to both sides
y=4. But the first way I did it is the best. Separate variables from numbers, then do what you need to in order to make the coefficient in front of the variable 1.